OpenStudy (anonymous):

What is the equation of a line that passes through the point (-2,4) and is parallel to the line y=3? A) x=-2 B) y=4 C) x=4 D) y=-2

OpenStudy (texaschic101):

y = 3 is a horizontal line with a 0 slope. A parallel line will have the same slope. now use y = mx + b slope(m) = 0 (-2,4) x = -2 and y = 4 now sub 4 = 0(-2) + b 4 = 0 + b 4 = b therefore your equation is : y = 0x + 4, or just y = 4

OpenStudy (anonymous):

So, B?

OpenStudy (texaschic101):

yes.. B